How to solve this next two term sequence《1,8,27,64》please help me.
Answer by Edwin McCravy(20056)

《1,8,27,64》
That's
《1×1×1=1³=1, 2×2×2=2³=8, 3×3×3=3³=27, 4×4×4=4³=64》
So to extend it use 5×5×5=5³=125, 6×6×6=6³=216, etc.